OpenStudy (anonymous):

If f(x)= x^3 + x^2 + x +1, find a number c that satisfies the conclusion of the Mean Value Theorem on the interval [0,4]. (This seems very straight forward, but I keep getting the wrong answer).

OpenStudy (solomonzelman):

I am going to assume you know the Mean Value Theorem. Okay?

OpenStudy (anonymous):

Correct assumption.

OpenStudy (solomonzelman):

You need to: 1) Find f(0) 2) Find f(4) 3) Find the average rate of change betwen (0,f(0)) and (4,f(4)) then we will find point(s) that have the same slope as the one you find in step 3./

OpenStudy (solomonzelman):

( first, do the 1st 3 steps )

OpenStudy (anonymous):

I did all that. Somehow I'm messing up.

OpenStudy (solomonzelman):

ok, what was your slope of the secant from x=0 to x=4?

OpenStudy (anonymous):

21

OpenStudy (solomonzelman):

Ok, f(0)=1 f(4)=85 slope of the secant = (85-1)/(4-0) = 84/4 = 21 Correct!

OpenStudy (solomonzelman):

Now, find f`(x)

OpenStudy (solomonzelman):

And then set the f`(x)=c, in this case 21, and we will see at which points does the function have an instanteneous slope of 21.

OpenStudy (anonymous):

I've done that twice. I must be making an error.

OpenStudy (solomonzelman):

you have a problem taking the derivative of ` x³+x²+x+1 ` , is that correct ?

OpenStudy (anonymous):

No. I get 3x^2 + 2x +1

OpenStudy (solomonzelman):

yes, that is correct. f`(x)=3x²+2x+1

OpenStudy (solomonzelman):

REMEMBER! The derivative of the function, is its (the function's) slope. Now, the slope of the secant in our case is 21. And this is why you need to set f`(x)=21, to see at which points will the derivative of the function (or the function's slope) be =21.

OpenStudy (anonymous):

Yes, I understand that. That's what I did. I think I must be making some stupid mistake, because I'm getting a different answer than the book.

OpenStudy (solomonzelman):

what is the answer in the book?

OpenStudy (solomonzelman):

and what answer have you previously been getting?

OpenStudy (anonymous):

I got x = -3 and x = 7/3. The book says it's (sqrt(61 - 1)/3.

OpenStudy (anonymous):

oops, -3 can't be right.

OpenStudy (solomonzelman):

yes, the book is correct.... :) (well, duh, but I actually did it...)

OpenStudy (solomonzelman):

when you set the equation 3x²+2x+1=21, you will get two solutions, and you will exclude one of them because this another solution is NOT on the interval of [0,4]. Can you solve 3x²+2x+1=21 ? (Shouldn't be a problem for you))

OpenStudy (solomonzelman):

you can use the quadratic formula (i would advise)

OpenStudy (anonymous):

You would think so, but apparently not.

OpenStudy (solomonzelman):

Ok, should I do this algebraic task for you?

OpenStudy (solomonzelman):

well, we are getting the calculus algorithm, so that wouldn't do much harm....

OpenStudy (anonymous):

Well, that seems to be where my brain is malfunctioning.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3x^2+2x+1=21 }\) \(\large\color{black}{ \displaystyle 3x^2+2x-20=0 }\) \(\large\color{black}{ \displaystyle \frac{-2\pm \sqrt{2^2-4(3)(-20)}}{2\cdot 3} }\) \(\large\color{black}{ \displaystyle \frac{-2\pm \sqrt{224}}{2\cdot 3} }\) \(\large\color{black}{ \displaystyle \frac{-2\pm \sqrt{61\cdot 4}}{2\cdot 3} }\) \(\large\color{black}{ \displaystyle \frac{-2\pm 2\sqrt{61}}{2\cdot 3} }\)

OpenStudy (solomonzelman):

from here you can proceed with no probs....

OpenStudy (anonymous):

OMG! Dyslexia strikes again. Instead of putting in 2^2, i put in 4 AND THEN squared that. Thanks so much!! This kind of thing makes me crazy.

OpenStudy (solomonzelman):

oh, it is just a minor mistake... messes everything up but happens to everybody:)

OpenStudy (anonymous):

How do you get the math characters in the display like that?

OpenStudy (solomonzelman):

I am using "latex"

OpenStudy (anonymous):

OK, thanks. I'll have to research that some time. Thanks again.

OpenStudy (solomonzelman):

you can look up a "legendary latex tutorial" online, and you should find a tutorial made by thomaster

OpenStudy (anonymous):

Will do. Thanks.

OpenStudy (solomonzelman):

If you want, we can do a polynomial approximation of √61, using f(x)=√x, for x=64.

OpenStudy (solomonzelman):

but that is not required I guess....

OpenStudy (solomonzelman):

so, your only answer (that is in the interval [0,4] ) = \(\large\color{black}{ \displaystyle \frac{1 }{3} (\sqrt{61}-1) }\)

OpenStudy (solomonzelman):

any questions ?

OpenStudy (solomonzelman):

So (extra part) you have a function: \(\large\color{black}{ \displaystyle f(x)=\sqrt{x} }\) \(\large\color{black}{ \displaystyle L(x)=f(a)+f`(a)(x-a) }\) this is a linear approximation of the function f(x) near the a point x=a. in this case our function is f(x)=√x, and a=64. (our f`(x) is going to be 1/(2√x), and f`(a)=f`(8)=1/(2√64) ...) \(\large\color{black}{ \displaystyle L(x)=\sqrt{64}+\frac{1}{2\sqrt{64}}(x-64) }\) \(\large\color{black}{ \displaystyle L(x)=8+\frac{1}{16}(x-64) }\) this is the tangent line to √x, at x=64 now, you can plug in 61 to get an approzimate value for √61 \(\large\color{black}{ \displaystyle \sqrt{61}\approx 8+\frac{-3}{16} }\) \(\large\color{black}{ \displaystyle \sqrt{61}\approx 7+\frac{13}{16} }\)

OpenStudy (solomonzelman):

But, by taylor approximation, you would have a closer approximation. (that is, drawing a tangent the the f(x) polynomial of nth degree thorugh the point x=a) \(\color{black}{ \displaystyle {\rm Approximation~of~f(x)}_{\rm~at~~x~=~a}~~ \\[1.3em] \displaystyle =f(a)+f`(a)(x-a)+\frac{f``(a)}{2!}(x-a)^2+\frac{f```(a)}{3!}(x-a)^2+ \\[1.3em] \displaystyle {\bf ....}~+~+\frac{f^{(n)}(a)}{n!}(x-a)^n }\)

OpenStudy (solomonzelman):

in any case... won't overwhelm more than what I did already